OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10? What is the slope of the line passing through the points (0, −5) and (4, 2)?

OpenStudy (anonymous):

Okay. 2x - 4y = 10 => Factor out 2 from the left hand side, i.e. 2(x - 2y) = 10 x - 2y = 5 => Set x - 2y = 5 in the slope intercept form: y = mx + c, where m = co-efficient of x = slope, then x - 2y = 5 → y = mx + c - 2y = - x + 5 → y = mx + c 2y = x - 5 → y = mx + c y = x/2 - 5/2 → y = mx + c => From what you have above, m = slope of the line: 2x - 4y = 10 is 1/2 Thats it!

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @jhtbsig

OpenStudy (anonymous):

so what're the answers @Tom_Boy_Rebel